OpenStudy (anonymous):

Find the radius of convergence and interval of convergence of the series \[\sum_{n=1}^{∞}\] 2^n(x-2)^n / (n+2)!

OpenStudy (anonymous):

i am getting that it converges no matter what.

OpenStudy (anonymous):

by ratio test. but i could be wrong so let me try again

OpenStudy (anonymous):

can u show me the steps u did?

OpenStudy (anonymous):

\[\frac{a_{n+1}}{a_n}=\frac{2(x-2)}{n+3}\]

OpenStudy (anonymous):

oh n+3

OpenStudy (anonymous):

i computed the ratio and that is what i got. \[a_{n+1}=\frac{2^{n+1}(x-2)^{n+1}}{(n+3)!}\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

so \[\frac{a_{n+1}}{a_n}=\frac{2^{n+1}(x-2)^{n+1}(n+2)!}{(n+3)!2^n(x-2)^n}\]

OpenStudy (anonymous):

after canceling my fool head off i got \[\frac{a_{n+1}}{a_n}=\frac{2(x-2)}{n+3}\]

OpenStudy (anonymous):

whose limit as n goes to infinity is 0 regardless of x. you should check my algebra though!

OpenStudy (anonymous):

yah its between 1/2 < x < 7/2 when it converges

OpenStudy (anonymous):

on the other hand n! grows way faster than 2^n so it is not surprising that this would converge

OpenStudy (anonymous):

really? then i made a mistake. i wonder what it was?

OpenStudy (anonymous):

what did you do?

OpenStudy (anonymous):

wait my bad its 0<x<3/2

OpenStudy (anonymous):

what did you do with this \[\frac{a_{n+1}}{a_n}=\frac{2(x-2)}{n+3}\]?

OpenStudy (anonymous):

yah that but i have to find the interval

OpenStudy (anonymous):

well it has been a while but as i recall you just have to make sure that for fixed x the the limit as n goes to infinity of this thing has to be less than 1 in absolute value. for a fixed x this limit is zero. so unless i forgot something important this converges for all x

OpenStudy (anonymous):

you will notice that there is no "n" in the numerator. so i think it converges for all x

OpenStudy (anonymous):

almost positive it does in fact.

OpenStudy (anonymous):

i certainly converges for x = 10 because i just tried it look here http://www.wolframalpha.com/input/?i=sum+2^n%2810-2%29^n+%2F+%28n%2B2%29!

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

usually you get something like \[\frac{nx}{2n}\] and then you have to make sure that\[|\frac{nx}{2n}|\leq1\] so \[-2 \leq x \leq 2\] is your radius. but in your example the limit is zero no matter what. just thought i would mention it

OpenStudy (anonymous):

thanks